OpenStudy (anonymous):

One pipe can fill a swimming pool in 5 hours less than another. Together they fill the swimming pool in 5 hours. How long would it take each pipe to fill the tank alone?

hartnn (hartnn):

let one of th pipe take 'x' hrs then other would take (x-5) hrs, right ? Now, do you know how to make an equation of this : 'Together they fill the swimming pool in 5 hours.' ??

OpenStudy (anonymous):

5=(x-5)+5

OpenStudy (anonymous):

so x =5

hartnn (hartnn):

nope, if A can do the work in y days , B can do the work in z days then they both together will do the work in (1/y+1/z) days so, here it'll be \((1/x)+1/(x-5)=5\) can you solve this ?

hartnn (hartnn):

sorry, its \((1/x)+1/(x-5)=1/5\)

OpenStudy (anonymous):

y is i 1/x? y division?

hartnn (hartnn):

ok, if A can do the work in 'y' hrs, then he will do (1/y) of the work in 1 hour , right ? got this ? (we take unit time, like 1 hr, 1 day, ....)

hartnn (hartnn):

one of the pipe take 'x' hrs to fill, then that pipe would fill 1/x of the pool in 1 hr.

OpenStudy (anonymous):

i see

hartnn (hartnn):

other would take (x-5) hrs, then other pipe would take 1/(x-5) of the pool in 1 hr.

OpenStudy (anonymous):

then u would solve for x?

hartnn (hartnn):

yup, solve for x in 1/x +1/(x-5) = 1/5

OpenStudy (anonymous):

\[(15+-\sqrt{125})/2\]

hartnn (hartnn):

yeah, i get the same weird answer :O

hartnn (hartnn):

but the equation 1/x +1/(x-5) =1/5 is absolutely correct.

OpenStudy (anonymous):

okay thnks

hartnn (hartnn):

welcome ^_^